OpenStudy (anonymous):

simplify 36/100, 53/100, 17/100, 24/100,

OpenStudy (mathstudent55):

53/100 and 17/100 are already simplified 36/100 = 9/25 24/100 = 6/25

OpenStudy (skullpatrol):

To simplify a fraction find numbers that divide evenly into the numerator and denominator.

OpenStudy (skullpatrol):

24/100 = 6/25 because 4 divides into 24; 6 times and 4 divides into 100; 25 times.